a coin collector has $45 in quarters and dimes. How many coins of each does he have is the total number of his coins is 240?
Thank you very much!

Answer by ptaylor(1894)

Let x=Number of dimes
Then 240-x=Number of quarters
Note: Whatever monetary units we deal in (pennies, nickels, dollars, etc.), we have to be consistent. I choose to deal in pennies.
Now we know that the amount of money in dimes (x)(10) plus the amount of money in quarters (240-x)(25)equals the total amount of money (4500). Thus, our equation to solve is:
10x+(240-x)(25)=4500 simplifying, we get:
10x+6000-25x=4500 or
-15x=-6000+4500=-1500
x=100 dimes
240-x=240-100=140 quarters
Hope this helps ----ptaylor